OpenStudy (anonymous):

Trains A and B are traveling in the same direction on parallel tracks. Train A is traveling at 80 miles per hour and train B is traveling at 96 miles per hour. Train A passes a station at 11:25 A.M. If train B passes the same station at 11:40 A.M., at what time will train B catch up to train A?

OpenStudy (anonymous):

help me please

OpenStudy (anonymous):

no

OpenStudy (anonymous):

jk, i will help

OpenStudy (anonymous):

basically what we can do here is set up a system of equations

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

now we will work with this equation: D=R*T

OpenStudy (anonymous):

Distance=Rate*Time

OpenStudy (anonymous):

Okay. so.....

OpenStudy (anonymous):

now, the distance traveled by A when B catches up is modeled by the equation: 80*T

OpenStudy (anonymous):

im not sure how to set it up

OpenStudy (anonymous):

for traveled distance?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now the distance traveled by B when it catches up is, remeber is i15 minutes behind, can be modeled by: 96*(T-0.15)

OpenStudy (anonymous):

now we have two equations: 80*T and 96*(T-0.15)

OpenStudy (anonymous):

0.15... so would one hour be 0.60

OpenStudy (anonymous):

i mean 96*(T-0.25) sorry

OpenStudy (anonymous):

so, we have two equations : 80*T and 96*(T-0.25)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

now we are looking for the point at which these two trains meet, to determine this we set both equations equal to each other, like so: :80*T =96*(T-0.25) that means: 80T=96T-24

OpenStudy (anonymous):

Now with this equation: 80T=96T-24 lets subtract 96T from both sides: -16T=-24, now divide both sides by -16 and get: T=(-24)/(-16)=(3)/(2)

OpenStudy (anonymous):

or 1.5 hours, that means the two trains meet 1hour and 30 min after train A passes the station

OpenStudy (anonymous):

you can thank people who answer you questions by clicking the good answer button